The substitution method has my head twisted can someone help me please???
-2y+ 6x=50
2y-11x=66

Answer by lwsshak3(6490)

The substitution method has my head twisted can someone help me please???
-2y+ 6x=50
2y-11x=66
**
solve for 2y in first equation
2y=6x-50
sub 2y into second equation
6x-50-11x=66
-5x=116
x=-116/5